# Constrained Portfolio Analysis in High Dimensions: Tracking Error and Weight Constraints

### v1

@@ 5 Tracking Error with Inequality Constraints on the Weights of the Portfolio
$\left|\frac{\hat{\boldsymbol{w}}_{est}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}_{est}}{\boldsymbol{w}_{est}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{est}^{*}}-1\right|=O_{p}(r_{w1}\bar{s}r_{T}K^{9/2})=o_{p}(1),$
but if $\kappa w_{u}>\omega$ (binding weight restriction),
$\left|\frac{\hat{\boldsymbol{w}}_{est}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}_{est}}{\boldsymbol{w}_{est}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{est}^{*}}-1\right|=O_{p}(r_{w2}\bar{s}^{2}r_{T}^{2}K^{6})=o_{p}(1).$
@@
$\left|\left[\frac{\widehat{SR}_{est}}{SR_{est}^{*}}\right]^{2}-1\right|=O_{p}(r_{w1}\bar{s}r_{T}K^{9/2})=o_{p}(1).$
otherwise if $\kappa w_{u}>\omega$ (binding weight restriction) then
$\left|\left[\frac{\widehat{SR}_{est}}{SR_{est}^{*}}\right]^{2}-1\right|=O_{p}(r_{w2}\bar{s}^{2}r_{T}^{2}K^{6})=o_{p}(1).$
Remarks.
1. $\epsilon$ is well defined as an upper bound that goes to zero. Specifically, $|\kappa(\hat{w}_{u}-w_{u})|\leq\epsilon$ wpa1, and given that $\kappa$ is a positive constant, we have $\epsilon\mathrel{\mathop{:}}=\bar{s}^{2}r_{T}l_{T}K^{7/2}\to 0$ by Assumption 7(i). This can be seen in Lemma A.2(iv).
2. So we exclude only a small-local to zero- neighborhood of zero in the indicator function $|\omega-\kappa w_{u}|>2\epsilon>0$, since $\epsilon\to 0$. So the binding versus nonbinding decisions are correct, as shown in the proof of Theorem 3, with probability approaching one (wpa1).
3. Our Theorem is new, and encompasses portfolio decision based on binding versus nonbinding constraints taken into account into consideration stochastically.
4. Main ingredient of the proof is when the constraint is binding, $\kappa w_{u}>\omega$, we have wpa1, $\kappa\hat{w}_{u}>\omega$, and the same is true for nonbinding constraint $\kappa w_{u}<\omega$ implies wpa1, $\kappa\hat{w}_{u}<\omega$. We can select binding versus nonbinding constraints correctly, with probability approaching one, and this decision is incorporated into our proofs.
@@ 6 Simulations
In this section we try to analyze our estimator in relation to Theorems, and also compare with existing methods. But we should emphasize that other methods do not have econometric-statistical theories for constrained high dimensional portfolio formation.
We generate the return data $\{y_{t}\}_{t=1}^{T}$ based on the following factor model with $K=3$ common factors. Specifically, we assume the factors follow the AR(1) processes:
$f_{i,t}=\mu_{f,i}+\alpha_{i}f_{i,t-1}+e_{i,t},\quad i=1,2,3$ (11)
@@
where $e_{i,t}\sim N(0,1-\alpha_{i}^{2})$, and $f_{i,0}=0$, for $i=1,2,3$. We set $\mu_{f,i}=0.005$ for $i=1,2,3$ and $\alpha_{1}=0.03,\alpha_{2}=-0.05,\alpha_{3}=-0.05$. For factor loadings $\boldsymbol{b}_{i}=\left(b_{i,1},b_{i,2},b_{i,3}\right)^{\top}$, $i=1,2,\cdots,p$, and the idiosyncratic errors $\boldsymbol{u}_{t}=(u_{1,t},\ldots,u_{p,t})^{\top}$, we draw their values from normal distributions. Specifically, $b_{i,j}\sim N(\mu_{b,j},1)$, where $\mu_{b,1}=-0.1,\mu_{b,2}=\mu_{b,3}=0.1$; and $\boldsymbol{u}_{t}\sim N(0,\Sigma_{u})$ with $\boldsymbol{\Sigma}_{u}$ has the Toeplitz form such that the $(i,j)$-th element of $\boldsymbol{\Sigma}_{u}$ is $0.25^{|i-j|}$. The population covariance of asset returns is $\boldsymbol{\Sigma}_{y}=\boldsymbol{B}\boldsymbol{B}^{\prime}+\boldsymbol{\Sigma}_{u}$, where $\boldsymbol{B}=(b_{1},\ldots,b_{p})^{\prime}$. Denote inverse of $\boldsymbol{\Sigma}_{y}$ as $\boldsymbol{\Theta}$. Set $\boldsymbol{c}\boldsymbol{o}\boldsymbol{v}\boldsymbol{f}_{t}=I_{3}$. We consider two $[p,T]$ combinations such that $p=0.8T$ and $p=1.2T$ with $T=100,150$. The benchmark market index $\boldsymbol{m}$ is an equal-weighted portfolio. We repeat the simulations $n_{sim}=100$ times.
We report the performances of 1), the “Oracle”, which is the portfolio given the population expected return $\boldsymbol{\mu}$ and risk aversion parameter $\Xi$ and plug in the analytical solution for $\boldsymbol{w}_{d}^{*}$ in (3), $\boldsymbol{w}_{cp}^{*}$ in (7), or $\boldsymbol{w}_{op}^{*}$ given the target constraints, namely, the TE only, TE with weight equality constraint, and TE with weight inequality constraint, respectively; 2), the portfolio with no constraint, “NCON”, which is the oracle portfolio without the aforementioned constraints; and 3) the benchmark, “Index”, which is the market index portfolio. We compare the performance of the proposed portfolio (“rb-NW”, stands for the plug-in $\hat{\boldsymbol{w}_{d}}$ or $\hat{\boldsymbol{w}_{cp}}$ or $\hat{\boldsymbol{w}}_{op}$ using the residual based nodewise in Caner et al. (2023)) with aforementioned benchmarks. We also compare rb-NW with other popular method for covariance/precision matrix estimation, specifically, the nodewise estimator Callot et al. (2021), the POET Fan et al. (2013), the (single factor) nonlinear shrinkage estimator Ledoit and Wolf (2017). These are later referred to as “comparison portfolios”. Similar to the proposed portfolio based on rb-NW, we construct these comparison portfolios using the plug-in covariance estimators from their respective methods.
To verify the main theoretical results and make comparisons of different portfolios, we report the following statistics, by presenting $\boldsymbol{w}_{1}^{*},\hat{\boldsymbol{w}}_{1}$ as representative portfolios in Theorems 1-3, and will be explained below:
$TE=\sqrt{(\hat{\boldsymbol{w}}_{1}-\boldsymbol{m})^{\prime}{\boldsymbol{\Sigma}_{y}}(\hat{\boldsymbol{w}}_{1}-\boldsymbol{m})},$
is the tracking error, where $\boldsymbol{m}$ is the benchmark portfolio, $\hat{\boldsymbol{w}}_{1}$ is the estimated weight, $\boldsymbol{\Sigma}_{y}$ is the covariance matrix of returns. The risk error is computed by
[…]
The Sharpe ratio error is defined by
[…]
where $\widehat{SR}=\hat{\boldsymbol{w}}_{1}^{\prime}\boldsymbol{\mu}/\sqrt{\hat{\boldsymbol{w}}_{1}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}_{1}}$, $SR={\boldsymbol{w}}_{1}^{*\prime}\boldsymbol{\mu}/\sqrt{{\boldsymbol{w}}_{1}^{*\prime}\boldsymbol{\Sigma}_{y}{\boldsymbol{w}}_{1}^{*}}$ and $\boldsymbol{\mu}$ is population mean of assets returns. The weight error is defined as
[…]
We report the averages of the $n_{sim}$ simulation repeats. Bold faced numbers show category (such as TE, weight, risk, SR) winners. To be more specific, for Tables 1 and 2 with TE constraint only, we use $\hat{\boldsymbol{w}}_{1}=\hat{\boldsymbol{w}}_{d}+\boldsymbol{m}$ and $\boldsymbol{w}_{1}^{*}=\boldsymbol{w}^{*}_{d}+\boldsymbol{m}$; for Tables 3 and 4 with TE and weight (equality) constraints, we use $\hat{\boldsymbol{w}}_{1}=\hat{\boldsymbol{w}}_{cp}+\boldsymbol{m}_{R}$ and $w_{1}^{*}=\boldsymbol{w}^{*}_{cp}+\boldsymbol{m}_{R}$. For Tables 5-6 with TE and weight inequality constraints, we use $\hat{\boldsymbol{w}}_{1}=\hat{\boldsymbol{w}}_{d}+\boldsymbol{m}$ if $\kappa\hat{w}_{u}<\omega$ and $\hat{\boldsymbol{w}}_{R}=\hat{\boldsymbol{w}}_{cp}+\boldsymbol{m}_{R}$ if $\kappa\hat{w}_{u}\geq\omega$. The theoretical $\boldsymbol{w}_{1}^{*}=\boldsymbol{w}^{*}_{d}+\boldsymbol{m}$ if $\kappa w_{u}<\omega$, and $\boldsymbol{w}_{1}^{*}=\boldsymbol{w}_{cp}^{*}+\boldsymbol{m}_{R}$ if $\kappa w_{u}\geq\omega$. For the TE constraints, we set the target TE to 0.1, 0.2, and 0.3, respectively; the weight restriction (if any) is imposed on the first $10$ stocks, that is, $1_{R}=(\underbrace{1,\dots,1}_{1\times 10},0,\ldots,0)^{\prime}$.
We show the best portfolio among the comparison portfolios (portfolios except Oracle, NCON, Index portfolios) with bold letter. The simulation results for TE constrained portfolios are shown in Tables 1-2. Recall that the reported statistics, weight, risk, SR, are the estimation errors deviating from the oracle values. First of all, among all strategies considered here, we can observe that rb-NW is the top performer in terms of SR estimation in Table 1. In Table 2, where $T=150$, rb-NW has the best performance in SR error for both $p<T$ and $p>T$, and all TE levels, except for $p=120$ and TE=0.3, where SFNL has slightly better SR error than rb-NW (0.044 compared to 0.047). It also has in general superior performance in other metrics such as weight and risk errors versus other methods. rb-NW also achieves desirable TE control (closest to the target TE levels) and weight error minimization. SFNL also has relatively good performance, specifically, it sometimes achieves lower weight error than that of rb-NW.
At last, the performance of rb-NW portfolio increase in general with the sample size. For example, when we fix the portfolio size at p=120, and increase the sample size from T=100 to 150, the errors of metrics that we consider, of rb-NW decrease in TE, weight, risk metrics, and in case of SR it decreases in all cases except $TE=0.3$ scenario.
Table 1: Simulation results for TE constraint, sample size $T=100$.
p=80, T=100 p=120, T=100
@@
TE Weight-ER Risk-ER SR-ER TE Weight-ER Risk-ER SR-ER
Oracle 0.100 0.000 0.000 0.000 0.102 0.000 0.000 0.000
NCON 0.255 0.987 0.863 4.030 0.212 0.975 0.802 2.531
Index $0$ 0.184 0.965 $0$ 0.261 0.963
NW 0.537 0.168 4.301 3.276 0.918 0.241 17.009 2.446
rb-NW 0.129 0.085 0.396 0.780 0.155 0.111 1.002 0.869
POET 0.278 0.118 1.539 2.175 0.303 0.163 3.613 1.341
NLS 0.165 0.105 0.757 1.417 0.178 0.154 1.953 1.233
SFNL 0.157 0.085 0.555 1.105 0.173 0.108 1.226 1.016
@@
TE Weight-ER Risk-ER SR-ER TE Weight-ER Risk-ER SR-ER
Oracle 0.199 0.000 0.000 0.000 0.203 0.000 0.000 0.000
NCON 0.301 0.973 0.623 1.189 0.268 0.950 0.523 0.737
Index $0$ 0.441 0.985 $0$ 0.553 0.982
NW 1.075 0.337 11.583 1.075 1.836 0.482 41.453 0.702
rb-NW 0.274 0.171 1.057 0.296 0.310 0.222 2.376 0.214
POET 0.550 0.236 4.087 0.781 0.606 0.326 8.256 0.327
NLS 0.331 0.210 2.034 0.542 0.355 0.308 4.657 0.303
SFNL 0.314 0.169 1.461 0.448 0.345 0.216 2.856 0.284
@@
TE Weight-ER Risk-ER SR-ER TE Weight-ER Risk-ER SR-ER
Oracle 0.297 0.000 0.000 0.000 0.307 0.000 0.000 0.000
NCON 0.368 0.960 0.439 0.563 0.347 0.924 0.341 0.338
Index $0$ 0.625 0.989 $0$ 0.729 0.986
NW 1.602 0.502 17.154 0.512 2.738 0.727 57.405 0.306
rb-NW 0.408 0.255 1.560 0.086 0.463 0.334 3.260 0.040
POET 0.820 0.351 6.025 0.345 0.904 0.491 11.180 0.108
NLS 0.493 0.314 3.011 0.210 0.529 0.464 6.402 0.057
SFNL 0.469 0.252 2.146 0.170 0.515 0.325 3.895 0.058
  • “Oracle” is the portfolio using the population $\mu$ and $\Xi$ and plug in the analytical solution for $w_{d}^{*}$ given the target TE constraint. “NCON” is the oracle portfolio (using the population $\mu$ and $\Xi$) without TE constraint. “Index” is the market index portfolio, where we consider equal weight for simplicity. “NW” is the portfolio using plug-in nodewise estimator (Callot et al., 2021). “rb-NW” is the proposed portfolio using residual based nodewise estimator for $\Xi$. “POET”,“ NLS ” “SFNL” are the portfolios using plug-in POET (Fan et al., 2013), nonlinear shrinkage and single factor nonlinear shrinkage (Ledoit and Wolf, 2017) estimators, respectively. “TE” means the tracking error. The other three statistics are the empirical version of the definitions in Theorems 1 and 2. Specifically, “Risk-ER”, “SR-ER”, “Weight-ER” are the Risk error, SR error, and Weight error defined in the main text. “-” means the statistic is not available.
Table 2: Simulation results for TE constraint, sample size $T=150$.
p=120, T=150 p=180, T=150
@@
TE Weight-ER Risk-ER SR-ER TE Weight-ER Risk-ER SR-ER
Oracle 0.102 0.000 0.000 0.000 0.103 0.000 0.000 0.000
NCON 0.212 0.975 0.802 2.531 0.229 0.984 0.838 2.522
Index $0$ 0.261 0.963 $0$ 0.261 0.918
NW 0.896 0.233 16.005 2.442 1.325 0.292 29.527 2.429
rb-NW 0.133 0.089 0.596 0.616 0.158 0.114 0.912 0.816
POET 0.291 0.136 2.695 1.142 0.415 0.350 7.574 1.345
NLS 0.157 0.131 1.205 0.947 0.168 0.188 2.015 1.101
SFNL 0.144 0.089 0.753 0.735 0.164 0.115 1.064 0.922
@@
TE Weight-ER Risk-ER SR-ER TE Weight-ER Risk-ER SR-ER
Oracle 0.203 0.000 0.000 0.000 0.207 0.000 0.000 0.000
NCON 0.268 0.950 0.523 0.737 0.278 0.968 0.592 0.814
Index $0$ 0.553 0.982 $0$ 0.534 0.958
NW 1.792 0.466 38.963 0.704 2.651 0.585 73.187 0.791
rb-NW 0.266 0.178 1.406 0.153 0.317 0.227 2.176 0.244
POET 0.581 0.272 6.025 0.275 0.831 0.701 18.802 0.353
NLS 0.314 0.262 2.867 0.221 0.336 0.377 4.892 0.267
SFNL 0.288 0.177 1.733 0.205 0.328 0.223 2.510 0.297

### v2

@@ 5 TE with Inequality Constraints on the Weights of the Portfolio
$\left|\frac{\hat{\boldsymbol{w}}_{est}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}_{est}}{\boldsymbol{w}_{est}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{est}^{*}}-1\right|=\left|\frac{\hat{\boldsymbol{w}}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}}{\boldsymbol{w}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}^{*}}-1\right|=O_{p}(r_{w1}K^{5/2})=O_{p}(\bar{s}l_{T}K^{4})=o_{p}(1),$
otherwise if $\kappa w_{u}>\omega$ (binding weight restriction), then
$\left|\frac{\hat{\boldsymbol{w}}_{est}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}_{est}}{\boldsymbol{w}_{est}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{est}^{*}}-1\right|=\left|\frac{\hat{\boldsymbol{w}}_{R}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}_{R}}{\boldsymbol{w}_{R}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{R}^{*}}-1\right|=O_{p}(r_{w1}K^{5/2})=O_{p}(\bar{s}l_{T}K^{4})=o_{p}(1).$
@@
$\left|\left[\frac{\widehat{SR}_{est}}{SR_{est}^{*}}\right]^{2}-1\right|=\left|\left[\frac{\widehat{SR}}{SR}\right]^{2}-1\right|=O_{p}(r_{w1}K^{5/2})=O_{p}(\bar{s}l_{T}K^{4})=o_{p}(1).$
otherwise if $\kappa w_{u}>\omega$ (binding weight restriction), then
$\left|\left[\frac{\widehat{SR}_{est}}{SR_{est}^{*}}\right]^{2}-1\right|=\left|\left[\frac{\widehat{SR}_{R}}{SR_{R}^{*}}\right]^{2}-1\right|=O_{p}(r_{w1}K^{5/2})=O_{p}(\bar{s}l_{T}K^{4})=o_{p}(1).$
Remarks. 1. $\epsilon$ is well defined as an upper bound that goes to zero. Specifically, $|\kappa(\hat{w}_{u}-w_{u})|\leq\epsilon$, with probability approaching 1 (wpa1), and given that $\kappa$ is a positive constant, we have $\epsilon:=\bar{s}l_{T}K^{3/2}\to 0$ by Assumption 8(i). This can be seen in Lemma A.2(iv).
2. We exclude only a small “local to zero” neighborhood of zero in the indicator function $|\omega-\kappa w_{u}|>2\epsilon>0$, since $\epsilon\to 0$.
3. Our theorem is new and encompasses portfolio decisions based on binding versus nonbinding constraints taken into account stochastically.
4. The main ingredient of the proof is when the constraint is binding, we have wpa1, $\kappa\hat{w}_{u}>\omega$; and the same is true for nonbinding constraint, that $\kappa\hat{w}_{u}<\omega$, wpa1. We can select binding versus nonbinding constraints correctly, with probability approaching 1, and this decision is incorporated into our proof of Theorem 3.
5. One question is whether inequality constrained portfolios have less estimation errors compared with an unconstrained portfolio in $l_{1}$ norm based weight estimation. The answer depends on whether the constraints are binding or not. If the constraints are binding, Remark 5 of Section 4 will be in effect. On the other hand, if the constraints are not binding, Remark 6 of Section 3 will be the answer.
6. One key issue is that short-sale constraints. These take the form of $w_{j}\geq 0$, $j=1,\cdots,p$. Short sale constraint is a floor constraint on weight. In Appendix A.8, we show that floor constraints on weights on assets in a portfolio can be used as cap constraints on weights on the remaining assets in the portfolio. In other words, if we want to put short-sales constraints on assets 1 and 5, these will be equivalent to cap constraints on all other assets (i.e., all assets except asset 1 and asset 5) in our portfolio. Our method can only work in two limited sub-cases but cannot cover a short sales constraint for each asset in a portfolio. First, our short-sale constraint can be applied to only one asset in our portfolio, say $j$ th asset, $w_{j}\geq 0$. The second possibility is not a classic short sales constraint but a short sales constraint on a group of assets jointly in a portfolio. This group of assets can have their sum of total weights to be larger than equal to zero. Appendix A.8 contains derivations of weights and more discussion about short sales constraints. Also, in empirics, we provide results on how much our portfolios have short positions.
7. One can note that all the rates in Theorems 1(i)-2(i)-3(i), and Theorems 1(ii)(iii), 2(ii)(iii), 3(ii)(iii) are the same. This is also true with Lemma A.3 and Lemma A.4. So even though there are different restrictions in the portfolios, such as TE and weight constraints, the rate of convergence is the same. The reason for this is estimation of unconstrained maximum Sharpe Ratio weights $\boldsymbol{\Theta}\boldsymbol{\mu}/\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}$ derives the rates. The detailed discussion is in Appendix B.4. The practical implication of this result may be that, regardless of restrictions, if the portfolio manager wants to achieve a target return, there will be no difficulty stemming from restrictions to achieve that target return.
@@ 6 Simulations
In this section, we provide a brief simulation study to illustrate the finite sample performance of CROWN. We only show simulations with TE constraint and certain $p,T$ combinations, and the rest of the exercise, with weight, TE+weight, and inequality constraints, with more $p,T$ combinations and DGPs are in Appendix. In Tables 3-4 we report estimation errors of the weight, risk, and Sharpe Ratio. We also provide estimated Sharpe Ratio (SR), and TE. All the estimation errors and estimated SR and TE are obtained by averaging across 200 iterations.
The excess return data $\{\boldsymbol{y}_{t}\}_{t=1}^{T}$ is generated from a standard three-factor model whose parameters are calibrated from the actual data. The excess asset return at time t, for j-th asset is:
$y_{j,t}=b_{j,1}f_{1,t}+b_{j,2}f_{2,t}+b_{j,3}f_{3,t}+u_{j,t},\quad j=1,...,p;\ t=1,...T.$ (11)
@@
Specifically, we assume the factors $\boldsymbol{f}_{t}=(f_{1,t},f_{2,t},f_{3,t})^{\prime}$ follow a stationary VAR(1) model, $\boldsymbol{f}_{t}=\boldsymbol{c}_{f}+\boldsymbol{\Pi}_{f}\boldsymbol{f}_{t-1}+\boldsymbol{e}_{t},$ with errors following a normal distribution with zero mean and covariance matrix $\boldsymbol{c}ov(\boldsymbol{e}_{t})=\boldsymbol{\Sigma}_{f}-\boldsymbol{\Pi}_{f}\boldsymbol{\Sigma}_{f}\boldsymbol{\Pi}_{f}^{\prime}$, which is proved in Appendix A.1. For factor loadings $\boldsymbol{b}_{j}=\left(b_{j,1},b_{j,2},b_{j,3}\right)^{\top}$, $j=1,2,\cdots,p$, and the idiosyncratic errors $\boldsymbol{u}_{t}=(u_{1,t},\ldots,u_{p,t})^{\top}$, we draw their values from normal distributions. Specifically, $\boldsymbol{b}_{j}\sim N_{3}(\boldsymbol{\mu}_{b},\boldsymbol{\Sigma}_{b})$; and $\boldsymbol{u}_{t}\sim N(0,\boldsymbol{\Sigma}_{u})$ with $\boldsymbol{\Sigma}_{u}$ has the Toeplitz form such that the $(i,j)$ th element of $\boldsymbol{\Sigma}_{u}$ is $0.25^{|i-j|}$. Factor loadings have the following mean and variance-covariance matrix in Table A.1.
Table 1: Parameters of $\boldsymbol{b}_{j}$
$\boldsymbol{\mu}_{b}$ $\boldsymbol{\Sigma}_{b}$
1.0166 0.0089 0.0013 0.0046
0.5799 0.0013 0.2188 -0.0134
0.2937 0.0046 -0.0134 0.1491
Then we set the parameters of the design in factors and errors as in Table A.2 with $\boldsymbol{\Sigma}_{f}$ as the covariance matrix of $\boldsymbol{f}_{t}$.
Table 2: Parameters of $\boldsymbol{f}_{t}$
$\boldsymbol{c}_{f}$ $\boldsymbol{\Sigma}_{f}$ $\boldsymbol{\Pi}_{f}$
0.0445 1.5016 0.1338 0.1682 -0.1204 0.1555 -0.0324
0.0060 0.1338 0.3667 -0.0310 -0.0074 -0.0378 0.00318
0.0021 0.1682 -0.0310 0.6017 -0.0027 0.0031 0.01669
Tables A.1-A.2 are obtained through the following procedure. We utilize the daily data on returns of 100 industrial portfolios, which are intersections of 10 portfolios formed on size and 10 portfolios formed on book-to-market ratio, and the Fama/French 3 factors from the website of Kenneth French, at https://mba.tuck.dartmouth.edu/pages/faculty/ken.french/. We use 20-year data from 2003-01-01 to 2022-12-31 for calibration. We use $p/T=2$ for $p=80,320$ on the left panels of Tables 3-4 respectively with $TE=0.1,0.2,0.3$. Then we use $p/T=0.8$, with $p=80,320$ on the right panels of Tables 3-4 respectively. The benchmark market index $\boldsymbol{m}$ is an equal-weighted portfolio.We also report two more benchmark portfolios: 1), the “Oracle”, which is the portfolio given the population expected return $\boldsymbol{\mu}$ and covariance $\boldsymbol{\Sigma}_{y}$ and plug in the analytical solution for CROWN, given the required constraints; 2), the portfolio with no constraint, “NCON”, which is defined in Appendix A.7. We also report other popular methods for covariance/precision matrix estimation, specifically, the nodewise estimator Callot et al. (2021) will be called NW, the POET of Fan et al. (2013), the nonlinear shrinkage estimator and its single factor version Ledoit and Wolf (2017) will be abbreviated as NLS, SFNL respectively. We construct these comparison portfolios using the covariance estimators from their respective methods. To verify the main theoretical results and make comparisons, we report the following statistics $TE=\sqrt{(\hat{\boldsymbol{w}}-\boldsymbol{m})^{\prime}{\boldsymbol{\Sigma}_{y}}(\hat{\boldsymbol{w}}-\boldsymbol{m})},$ is the TE estimate, where $\hat{\boldsymbol{w}}$, $\boldsymbol{m}$ are the estimated weight and weight of the benchmark portfolio, in Theorem 1, respectively. The other metrics and their estimates, risk, Sharpe Ratio, weights are defined as in Theorem 1.
Bold-faced numbers in Tables 3-4 are the category winners among CROWN, NW, POET, NLS, and SFNL methods. First, in high dimensional Table 4, CROWN is the winner in all metrics. To give an example, at $p=320,T=160$ CROWN has SR estimation error of 0.0003 at TE=0.1, and the others range between 0.0059-0.0814. Then, as a second observation, in both Tables 3-4, on the left and right panels, when we increase $p,T$ fourfold but keep $p/T$ same, CROWN gets lower estimation rates, in line with theory. To give an example, with $p=80,T=40$ and with $TE=0.2$, CROWN has an estimation error of risk at 0.1622 in Table 3, and this error decreases in Table 4, with $p=320,T=160$, to 0.0015.
Table 3: Simulations with TE constraint, $p=80$
p=80, T=40 p=80, T=100
@@
TE Weight-ER Risk-ER SR-ER SR TE Weight-ER Risk-ER SR-ER SR
Oracle 0.1000 0.0000 0.0000 0.0000 0.0350 0.1000 0.0000 0.0000 0.0000 0.0350
NCON 0.4707 2.0415 0.1921 0.0248 0.0352 0.3698 1.1704 0.2394 0.0665 0.0361
Index - 0.4698 0.1665 0.0570 0.0339 - 0.4698 0.1697 0.0255 0.0345
NW 0.4013 1.7347 11.5532 0.0150 0.0346 0.2844 0.8272 0.2427 0.0337 0.0344
CROWN 0.1840 0.8981 0.0717 0.0055 0.0349 0.0984 0.1608 0.0027 0.0005 0.0349
POET 0.1192 0.8844 0.1316 0.0148 0.0347 0.1110 0.4889 0.0444 0.0046 0.0349
NLS 0.0961 0.6810 0.1056 0.0121 0.0347 0.1006 0.5363 0.0518 0.0049 0.0349
SFNL 0.1158 0.5463 0.0842 0.0113 0.0348 0.1077 0.4194 0.0404 0.0049 0.0349
@@
TE Weight-ER Risk-ER SR-ER SR TE Weight-ER Risk-ER SR-ER SR
Oracle 0.2000 0.0000 0.0000 0.0000 0.0350 0.2000 0.0000 0.0000 0.0000 0.0350
NCON 0.4707 2.2349 0.2741 0.0255 0.0352 0.3698 1.4176 0.3165 0.0655 0.0361
Index - 0.9396 0.0483 0.0579 0.0339 - 0.9396 0.0511 0.0264 0.0345
NW 0.4378 2.2713 12.4291 0.0166 0.0346 0.5688 1.6544 0.5052 0.0720 0.0337
CROWN 0.3680 1.7961 0.1622 0.0293 0.0345 0.1967 0.3216 0.0048 0.0012 0.0349
POET 0.2383 1.7688 0.2771 0.0407 0.0342 0.2220 0.9778 0.0919 0.0119 0.0348
NLS 0.1922 1.3621 0.2188 0.0319 0.0344 0.2012 1.0727 0.1069 0.0145 0.0347
SFNL 0.2315 1.0927 0.1726 0.0282 0.0345 0.2155 0.8389 0.0824 0.0131 0.0347
@@
TE Weight-ER Risk-ER SR-ER SR TE Weight-ER Risk-ER SR-ER SR
Oracle 0.3000 0.0000 0.0000 0.0000 0.0349 0.3000 0.0000 0.0000 0.0000 0.0349
NCON 0.4707 2.5137 0.3459 0.0265 0.0352 0.3698 1.7773 0.3841 0.0675 0.0361
Index - 1.4094 0.0555 0.0037 0.0339 - 1.4094 0.0529 0.0246 0.0345
NW 0.4816 2.4137 14.9933 0.0176 0.0346 0.8533 2.4817 0.7762 0.1068 0.0330
CROWN 0.5520 2.6942 0.2641 0.0622 0.0338 0.2951 0.4824 0.0066 0.0021 0.0349
POET 0.3575 2.6533 0.4294 0.0701 0.0337 0.3330 1.4667 0.1407 0.0223 0.0346
NLS 0.2883 2.0431 0.3349 0.0546 0.0340 0.3018 1.6090 0.1632 0.0273 0.0345
SFNL 0.3473 1.6390 0.2622 0.0472 0.0341 0.3232 1.2583 0.1246 0.0233 0.0345
  • “TE” is the Tracking Error. “Risk-ER”, “SR-ER”, “Weight-ER” are the estimation errors (defined in Appendix) of risk, SR and portfolio weights, respectively, as compared to the oracle. “SR” is the Sharpe Ratio. Due to space limitation we report the estimated return, risk in Appendix. “-” means the statistic is not available.
Table 4: Simulations with TE constraint, $p=320$
p=320, T=160 p=320, T=400
@@
TE Weight-ER Risk-ER SR-ER SR TE Weight-ER Risk-ER SR-ER SR
Oracle 0.1000 0.0000 0.0000 0.0000 0.0353 0.1000 0.0000 0.0000 0.0000 0.0353
NCON 0.3988 1.3717 0.3109 0.0705 0.0365 0.3975 1.3825 0.3109 0.0708 0.0365
Index - 0.6366 0.1593 0.0699 0.0340 - 0.6366 0.1624 0.0400 0.0346
NW 0.6485 2.4803 0.8161 0.0814 0.0338 0.6661 2.6621 0.8475 0.0689 0.0340
CROWN 0.0996 0.2142 0.0009 0.0003 0.0353 0.1001 0.1570 0.0008 0.0002 0.0353
POET 0.1752 5.1018 3.4305 0.0347 0.0345 0.1427 2.8335 1.8423 0.0199 0.0349
NLS 0.1088 2.7118 0.2809 0.0092 0.0351 0.1209 2.1571 0.1368 0.0043 0.0352
SFNL 0.1414 1.5372 0.1697 0.0059 0.0352 0.1236 1.0816 0.0797 0.0025 0.0352
@@
TE Weight-ER Risk-ER SR-ER SR TE Weight-ER Risk-ER SR-ER SR
Oracle 0.2000 0.0000 0.0000 0.0000 0.0354 0.2000 0.0000 0.0000 0.0000 0.0354
NCON 0.3988 1.8302 0.3878 0.0650 0.0365 0.3975 1.8374 0.3877 0.0653 0.0365
Index - 1.2733 0.0300 0.0747 0.0340 - 1.2733 0.0327 0.0450 0.0346
NW 1.2969 4.9607 1.8131 0.1476 0.0326 1.3322 5.3243 1.8713 0.1241 0.0331
CROWN 0.1992 0.4284 0.0015 0.0008 0.0353 0.1999 0.3140 0.0014 0.0005 0.0353
POET 0.2190 6.1584 5.2542 0.0386 0.0344 0.2039 3.5276 3.5383 0.0221 0.0349
NLS 0.2176 5.4237 0.5927 0.0347 0.0347 0.2417 4.3142 0.2847 0.0215 0.0350
SFNL 0.2829 3.0745 0.3447 0.0180 0.0350 0.2472 2.1633 0.1594 0.0086 0.0352